# High-order van Hove singularities and nematic instability in a kagome superconductor CsTi3Bi5

### v1

@@ Abstract
ATi3Bi5 (A = Cs or Rb) are emerging topological kagome metals that exhibit superconductivity and nematicity without intertwining translational symmetry-breaking charge orders. In this work, we explore the Fermionology of their titanium kagome electrons and identify a set of sublattice-pure, high-order van Hove singularities (VHSs) that can suppress charge ordering and enhance electronic correlations and superconductivity. Our calculations of charge susceptibility for kagome bands with both normal and high-order VHSs emphasize the role of these VHSs in driving electronic nematicity in CsTi3Bi5. Additionally, we compute the phonon spectrum and electron-phonon interactions for CsTi3Bi5 under pristine, doped, and kagome-exposed surface conditions, revealing its robustness against structural instabilities while enhancing the superconducting transition temperature. Our work positions ATi3Bi5 as a key platform for investigating superconductivity and electronic nematicity without translational symmetry-breaking states in kagome metals.
@@ I Introduction
Kagome materials, with uniformly interlocked hexagonal and triangular atomic networks, provide a fundamental platform for studying the interplay between frustrated geometry, non-trivial topology, and electronic correlations [1, 2]. Their kagome lattice results in Dirac cones that encode non-trivial topology, and van Hove singularities (VHSs) and flat bands, which drive electronic corrections. These electronic features have been observed in various kagome materials, such as the presence of Weyl fermions in the ferromagnet Co3Sn2S2 [3, 4], chiral spin textures and a significant anomalous Hall effect in the non-collinear antiferromagnet Mn3X (X = Sn and Ge)[5, 6], Dirac fermions and flat bands in FeSn and CoSn [7, 8], coexisting magnetism and superconductivity in Cu3Sn2 [9], and charge-density-wave (CDW) in antiferromagnetic FeGe [10].
The vanadium-based nonmagnetic kagome superconductors AV3Sb5 (A = K, Rb, or Cs) have realized a wide range of correlated states, including CDW [11, 12, 13, 14], pair density wave [15, 16], superconductivity [17, 18], and nematic order [19], often found in high-temperature superconductors [20]. The interplay between non-trivial topology and the correlated states in these materials has led to time-reversal symmetry-breaking phenomena [21, 22] and the potential emergence of Majorana zero modes [23]. Despite the diversity of quantum states in AV3Sb5, the mechanisms driving these states remain unresolved. This is due in part to the complex interplay of electronic and lattice degrees of freedom, both of which can drive the system into low-energy ordered states. In AV3Sb5, multiple VHSs near the Fermi level, arising from kagome electrons, create an environment prone to electronic instabilities, while the presence of soft phonon modes points to structural instabilities. For instance, the CDW order may result from electronic nesting at the VHSs or through electron-phonon interactions. Similarly, nematic order, which breaks rotational symmetry, might arise from electronic instabilities or as a result of translational symmetry breaking associated with CDW formation. Given these complexities, recent research has increasingly focused on exploring alternative kagome materials that are devoid of instabilities stemming from both electronic and phonon degrees of freedom.
@@
In this paper, we provide a comprehensive analysis of the electronic structure and electron-phonon interactions in CsTi3Bi5 using a uniform first-principles framework. We elucidate the multi-orbital nature of the Fermi surface, highlighting contributions from both Ti $d$ and Bi $p$ states, which hybridize to form multi-orbital Fermi pockets. The kagome bands near the Fermi level arise from Ti $d_{x^{2}-y^{2}}/d_{xy}$ states, with flat bands situated below the Fermi level, while VHSs and Dirac bands appear above it. Notably, both the VHSs are high-order and sublattice-pure, differing from those in AV3Sb5. Our detailed charge susceptibility calculations indicate that these high-order VHSs reduce Fermi surface nesting and can drive electronic instabilities. We also demonstrate the stability of CsTi3Bi5 against soft phonon modes and translational symmetry-breaking through analyses of phonon dispersions of pristine, charge-doped, and kagome-exposed surface conditions. Additionally, we explore strategies to enhance the superconducting Tc by tuning the Fermi level to align with flat bands or VHSs via hole and electron doping. Our results highlight that CsTi3Bi5 is a unique kagome material that possesses the necessary ingredients for realizing electronic instabilities such as nematicity and superconductivity, consistent with experimental findings.
@@ II Methodology
The superconducting $T_{c}$ of CsTi3Bi5 was obtained by evaluating the electron-phonon interaction based on the Eliashberg spectral function ($\alpha^{2}F(\omega)$) with Quantum Espresso [43] and EPW [44] codes. The energy spectrum in quantum espresso was calculated with PBE-GGA density functional and norm-conserving pseudopotentials [45] using the relaxed parameters obtained in VASP. A plane-wave energy cutoff of 60 Ry and a charge density cutoff of 480 Ry were used. The phonon spectrum was obtained through density functional perturbation theory. Notably, we confirmed the consistency of the electron and phonon dispersions calculated in VASP and quantum espresso codes and found an excellent agreement between them. For electron-phonon coupling, a fine 32$\times$32$\times$24 $k$-grid and 16$\times$16$\times$12 $q$-grid were used in the bulk Brillouin zone. The effects of additional carriers in CsTi3Bi5 were simulated by adding and removing electrons in the unit cell with a neutralizing uniform background charge. For the calculation of susceptibilities, a fine $200\times 200$ $k$ mesh was used at temperature $T=0.01$ K.
@@ III.1 Crystal structure and electronic state
Figure 1(b) presents the electronic structure of CsTi3Bi5 with spin-orbit coupling. The band structure features four dispersive band crossings at the Fermi level with similar characteristics at the $k_{z}=0$ and $k_{z}=\pi/c$ planes, which highlights the strong 2D nature of the Fermi pockets (Fig. 1(c)). Despite its complexity, several Kagome features are apparent in the partial density of states and the orbital-resolved band structure (Fig. 1(b)). The saddle-point VHSs (VHS1 and VHS2) are located above the Fermi level, while flat bands are found below it. The Dirac bands at the $K$ and $H$ points exhibit small gaps due to spin-orbit coupling. Importantly, the VHSs at the $M$ point are situated above the Fermi level compared to those in AV3Sb5 compounds due to the one-electron deficiency in Ti relative to V. The peaks in the density of states prominently reflect the presence of both VHSs and flat bands, with a strong contribution from Ti $d_{x^{2}-y^{2}}$/$d_{xy}$ states. Figure 1(c) illustrates the top view of the Fermi surface, comprising five 2D Fermi pockets: three centered at the $\Gamma$ point (one circular, one hexagonal, and one snowflake), one diamond-like pocket at the $M$ point, and one triangular pocket at the $K$ point. These pockets exhibit substantial hybridization between Ti $d$ and Bi $p$ bands, with varying character across the Fermi surface, as shown in the orbital-resolved Fermi band contours in Fig. 1(d). The central circular pocket is derived from Bi $p$ orbitals, while the snowflake pocket is primarily contributed by Ti $d_{x^{2}-y^{2}}/d_{xy}$ orbitals that form kagome bands. The remaining three pockets demonstrate strong $d-p$ hybridization, with their character alternating across the Fermi surface. This multi-orbital Fermi surface is similar to that of the FeSe superconductor, which also exhibits nematic instability. These Fermi surface characteristics are consistent with ARPES studies of CsTi3Bi5, which demonstrate polarization-selective band structure [28].
We emphasize that, despite the metallic nature of CsTi3Bi5, the band structure exhibits a local energy gap at all time-reversal invariant momentum points. These energy gaps enable the calculations of the $\mathbb{Z}_{2}$ invariant using the parity criterion [46]. Figure 1(e) presents the product of occupied band parities at various momentum points, along with the calculated nontrivial $\mathbb{Z}_{2}$ invariant. These results indicate that CsTi3Bi5 can host topological surface states within the local energy gaps, as reported in experimental studies [28, 26].
@@ III.2 VHSs and nematic instability
We now examine the kagome electronic features and their role in driving electronic nematic order in CsTi3Bi5. Figure 2(a) presents the Ti $d_{x^{2}-y^{2}}$/$d_{xy}$ orbital-resolved band structure, highlighting the Kagome electronic features near the Fermi level. The two VHSs, VHS1 and VHS2, are located at energies 0.22 eV and 0.74 eV above the Fermi level at $M$ point, respectively, while a flat band lies around -0.4 eV below it. These kagome bands cross the Fermi level, forming the Dirac cone at the $K$ point. To characterize the flavor of VHSs, we isolate the sublattice projections of the Bloch wave functions at VHSs, as shown in Fig.  2(b). Both VHS1 and VHS2 are sublattice pure, with each sublattice contributing VHS at different M points. Typically, in a kagome band structure, two VHSs exhibit a mix of sublattice pure and mixed characters; however, both VHSs here are sublattice pure without exhibiting mixed characters. This behavior arises from the three-dimensionality of VHS2, resulting from the mixing of Ti $d_{x^{2}-y^{2}}$/$d_{xy}$ with Ti $d_{z^{2}}$ orbitals. The sublattice-pure decoration of VHSs splits the three prominent particle-hole scattering channels in the hexagonal Fermi surface into six, thereby weakening the instabilities like CDW while promoting sublattice interference and long-range interactions [47].
To analyze the order of VHSs, we fit the energy dispersion around the saddle points (VHS1 and VHS2) along the $M-K-M$ direction using both quadratic and quartic polynomials, as shown in Fig. 2(c). The band dispersion is better described by a quartic polynomial $E=a+bk_{x}^{2}+ck_{x}^{4}$, with fitting parameters $a=0.22$, $b=1.07$, and $c=7.63$ for VHS1, and $a=0.74$, $b=-0.33$, and $c=-13.87$ for VHS2. The deviation of energy dispersion from a normal quadratic form indicates the presence of high-order VHSs (Fig. 2(e)). The high-order VHSs warp the ideal hexagonal Fermi surface associated with normal VHSs, reducing Fermi surface nesting and associated instabilities. Simultaneously, they enhance the DOS, thereby amplifying correlations-driven many-body effects. The sublattice-pure high-order VHSs in CsTi3Bi5 thus suggest a decreased likelihood of CDW instabilities while enhancing sublattice interference and electron correlation effects [48, 49, 50]. This interplay can ultimately strengthen electronic instabilities. Such phenomena have been observed in materials like CsV3Sb5 [51] and high $T_{c}$ cuprates [52], where significant many-body effects prevail.
@@
The aforementioned discussion indicates the possibility of electronic nematic instability in CsTi3Bi5, as observed in scanning tunneling microscopy measurements [27]. Notably, the nematic order in CsTi3Bi5 reduces the $C_{6}$ rotation symmetry to $C_{2}$ symmetry. Due to the sublattice pure VHSs, a simple onsite potential can break rotation symmetry, distorting the Fermi surface and lowering the density of states at VHSs. To test this phenomenology, we introduce an onsite potential in one of the sublattices, which effectively modifies the states in one of the three inequivalent $M$ points. The resulting $\chi({\bf q})$, shown in Fig. 2(h), demonstrates that the $C_{6}$ symmetry of the system is broken, leading to the emergence of $C_{2}$ symmetry. While this model captures the symmetry lowering in the absence of CDW ordering, the multi-orbital Fermi surface and sublattice interference result in non-uniform symmetry breaking, as seen in quasiparticle interference patterns, which warrant further investigation [27, 28, 29].
@@ III.3 Robustness of structural stability
We now examine the possibility of lattice instabilities in CsTi3Bi5. While the VHSs suggest possible electronic instabilities, they are situated slightly above the Fermi level. Therefore, it is essential to investigate dynamical instabilities that may arise from shifting the Fermi level to the VHSs or flat band energies. Figure 3(a) shows the phonon dispersion for pristine CsTi3Bi5, along with its electron and hole doping cases. In all cases, the phonon eigenvalues remain positive without any imaginary values across the full Brillouin zone. There are only minor changes in the phonon bands even under substantial doping of 1 electron (hole) per unit cell, which moves the Fermi level to the VHS1 (near flat band) energies. We also assessed the lattice stability by applying 1% compressive and tensile strains (Fig. 3(b)), which similarly yield stable phonon modes. The absence of imaginary phonon frequencies under various conditions suggests that CsTi3Bi5 is robust against lattice instabilities. Notably, even moving the Fermi level to the VHSs or flat bands through doping has no detrimental impact on lattice stability. These results align well with recent experimental reports indicating that CsTi3Bi5 exhibits no structural instabilities even when the VHSs are moved to the Fermi level through Cs doping [30].
@@
In bulk kagome materials, the kagome atomic layers are typically buried deep within the structure, complicating the assessment of their influence on structural instabilities. To examine the role of the Ti kagome layer, we calculated the surface phonon dispersions of CsTi3Bi5 for two surface terminations: one that directly exposes the Ti kagome layer and another that positions it just beneath the surface. The phonon spectra for both terminations, presented in Figs. 3(c) and 3(d), reveal no imaginary phonon frequencies in the surface Brillouin zone. These results indicate that CsTi3Bi5 retains structural stability even with the kagome layer is exposed at the surface, suggesting that any potential instabilities here would be driven by electronic factors. These characteristics position CsTi3Bi5 as a unique kagome system, especially in contrast to isostructural AV3Sb5 materials and bi-kagome layer systems such as AV6Sn6, which exhibit significant lattice instabilities in their surface-exposed kagome layers.

### v3

# High-order van Hove singularities and nematic instability in the kagome superconductor CsTi3Bi5

@@ Abstract
ATi3Bi5 (A = Cs or Rb) are emerging topological kagome metals that exhibit superconductivity and nematicity without intertwining translational symmetry-breaking charge orders. In this work, we explore the fermiology of their titanium kagome electrons and identify a set of sublattice-pure, high-order van Hove singularities (VHSs) that can suppress charge ordering and enhance electronic correlations and superconductivity. Our calculations of charge susceptibility for kagome bands with both normal and high-order VHSs emphasize the role of these VHSs in driving electronic nematicity in CsTi3Bi5. Additionally, we compute the phonon spectrum and electron-phonon interactions for CsTi3Bi5 under pristine, doped, and kagome-exposed surface conditions, revealing its robustness against structural instabilities while enhancing the superconducting transition temperature. Our work positions ATi3Bi5 as a key platform for investigating superconductivity and electronic nematicity without translational symmetry-breaking states in kagome metals.
@@ I Introduction
Kagome materials, with uniformly interlocked hexagonal and triangular atomic networks, provide a fundamental platform for studying the interplay between frustrated geometry, non-trivial topology, and electronic correlations [1, 2]. Their kagome lattice results in Dirac cones that encode non-trivial topology, and van Hove singularities (VHSs) and flat bands, which drive electronic correlations. These electronic features have been observed in various kagome materials, such as the presence of Weyl fermions in the ferromagnet Co3Sn2S2 [3, 4], chiral spin textures and a significant anomalous Hall effect in the non-collinear antiferromagnet Mn3X (X = Sn and Ge) [5, 6], Dirac fermions and flat bands in FeSn and CoSn [7, 8], coexisting magnetism and superconductivity in Cu3Sn2 [9], and charge-density-wave (CDW) in antiferromagnetic FeGe [10].
Among the various known kagome materials, the vanadium-based nonmagnetic kagome superconductors AV3Sb5 (A = K, Rb, or Cs) have realized a wide range of correlated states, including CDW [11, 12, 13, 14], chiral excitonic order [15], pair density wave [16, 17], superconductivity [18, 19], and nematic order [20], often found in high-temperature superconductors [21]. The interplay between non-trivial topology and the correlated states in these materials has led to time-reversal symmetry-breaking phenomena [22, 23] and the potential emergence of Majorana zero modes [24]. Despite the diversity of quantum states in AV3Sb5, the mechanisms driving these states remain unresolved. This is due in part to the complex interplay of electronic and lattice degrees of freedom, both of which can drive the system into low-energy ordered states. In AV3Sb5, multiple VHSs near the Fermi level, arising from kagome electrons, create an environment prone to electronic instabilities, while the presence of soft phonon modes points to structural instabilities. For instance, the CDW order may result from electronic nesting at the VHSs or through electron-phonon interactions. Similarly, nematic order, which breaks rotational symmetry, might arise from electronic instabilities or as a result of translational symmetry breaking associated with CDW formation. In addition, materials with HfFe6Ge6-type double-kagome lattice have been identified as a promising system to study the interplay between CDW and other electronic orders, as these materials exhibit not only CDW [25, 26, 27, 28, 29] but also nematicity [30], magnetism [31, 32], topological properties [33, 34] and superconductivity [35]. One notable example, ScV6Sn6, demonstrates the coexistence of CDW and nematicity, with the CDW primarily attributed to the Sc and Sn atoms, while the nematic order is largely associated with the kagome lattice. Despite exhibiting a $\sqrt{3}\times\sqrt{3}$ charge order, ScV6Sn6 maintains intra-unit-cell nematic order, preserving translational symmetry [30]. Given these complexities, recent research has increasingly focused on exploring alternative kagome materials that are devoid of instabilities stemming from both electronic and phonon degrees of freedom.
@@
In this paper, we provide a comprehensive analysis of the electronic structure and electron-phonon interactions in CsTi3Bi5 using a uniform first-principles framework. We elucidate the multi-orbital nature of the Fermi surface, highlighting contributions from both Ti $d$ and Bi $p$ states, which hybridize to form multi-orbital Fermi pockets. The kagome bands near the Fermi level arise from Ti $d_{x^{2}-y^{2}}/d_{xy}$ states, with flat bands situated below the Fermi level, while VHSs and Dirac bands appear above it. Notably, both the VHSs are high-order and sublattice-pure, differing from those in AV3Sb5. Our detailed charge susceptibility calculations indicate that these high-order VHSs reduce Fermi surface nesting and can drive electronic instabilities. We also demonstrate the stability of CsTi3Bi5 against soft phonon modes and translational symmetry-breaking through analyses of phonon dispersions of pristine, charge-doped, and kagome-exposed surface conditions. Moreover, we explore strategies to enhance the superconducting Tc by tuning the Fermi level to align with flat bands or VHSs via hole and electron doping. Our results highlight that CsTi3Bi5 is a unique kagome material that possesses the necessary properties for realizing electronic instabilities such as nematicity and superconductivity, consistent with experimental findings.
@@ II Methodology
The superconducting $T_{c}$ of CsTi3Bi5 was obtained by evaluating the electron-phonon interaction based on the Eliashberg spectral function ($\alpha^{2}F(\omega)$) with Quantum ESPRESSO (QE) [55] and EPW [56] codes. The energy spectrum in QE was calculated with PBE-GGA density functional and norm-conserving PBE pseudopotentials with standard accuracy from the PseudoDojo library [57] using the relaxed parameters obtained in VASP. A plane-wave energy cutoff of 60 Ry and a charge density cutoff of 480 Ry were used. The phonon spectrum was obtained through density functional perturbation theory. Notably, we confirmed the consistency of the electron and phonon dispersions calculated in VASP and QE codes and found a good agreement between them (see the Supplemental Materials (SM)) [58]. For electron-phonon coupling, a carefully converged fine 32$\times$32$\times$24 $k$-grid and 16$\times$16$\times$12 $q$-grid were used in the bulk Brillouin zone [58]. The effects of additional carriers in CsTi3Bi5 were simulated by adding and removing electrons in the unit cell with a neutralizing uniform background charge. For the calculation of susceptibilities, a fine $200\times 200$ $k$ mesh was used at temperature $T=0.01$ K.
@@ III.1 Crystal structure and electronic state
Figure 1(b) presents the electronic structure of CsTi3Bi5 with spin-orbit coupling. The band structure features four dispersive band crossings at the Fermi level with similar characteristics at the $k_{z}=0$ and $k_{z}=\pi/c$ planes, which highlights the strong 2D nature of the Fermi pockets (Fig. 1(c)). Despite its complexity, key Kagome features are apparent in the orbital-resolved band structure and partial density of states (Figs. 1(b) and 1(e)). The kagome bands are predominantly composed of Ti $d_{x^{2}-y^{2}}$/$d_{xy}$ orbitals with the saddle-point VHSs (VHS1 and VHS2) and Dirac points located above the Fermi level and flat bands below it. The Dirac points at the $K$ and $H$ points exhibit small gaps due to spin-orbit coupling. Notably, the VHSs and Dirac points in CsTi3Bi5 are located above the Fermi level, in contrast to those in AV3Sb5 materials, due to the one-electron deficiency of Ti compared to V. The position of the flat bands below the Fermi level is ascertained by the orbitals contributing to the kagome bands [59]. When the kagome bands are formed by $d_{xz}$/$d_{yz}$ orbitals, the flat band lies above the Dirac point. However, when the bands consist primarily of $d_{x^{2}-y^{2}}$/$d_{xy}$ orbitals, a sign change in the off-diagonal matrix elements shifts the flat band below the Dirac point, as seen in CsTi3Bi5.
Figure 1(c) illustrates the top view of the Fermi surface, comprising five 2D Fermi pockets: three centered at the $\Gamma$ point (one circular, one hexagonal, and one snowflake), one diamond-like pocket at the $M$ point, and one triangular pocket at the $K$ point. These pockets exhibit substantial hybridization between Ti $d$ and Bi $p$ bands, with varying character across the Fermi surface, as shown in the orbital-resolved Fermi band contours in Fig. 1(d). The central circular pocket is derived from Bi $p$ orbitals, while the snowflake pocket is primarily contributed by Ti $d_{x^{2}-y^{2}}/d_{xy}$ orbitals that form kagome bands. The remaining three pockets demonstrate strong $d-p$ hybridization, with their character alternating across the Fermi surface, resolving a momentum-dependent hybridization (see Figs. 1(d)-1(f)). This multi-orbital Fermi surface is similar to that of the FeSe superconductor, which also exhibits nematic instability. These Fermi surface characteristics are consistent with ARPES studies of CsTi3Bi5, which demonstrate polarization-selective band structure [40].
We emphasize that, despite the metallic nature of CsTi3Bi5, the band structure exhibits a local energy gap at all time-reversal invariant momentum points. These energy gaps enable the calculations of the $\mathbb{Z}_{2}$ invariant using the parity criterion [60]. Figure 1(g) presents the product of occupied band parities at various momentum points, along with the calculated nontrivial $\mathbb{Z}_{2}$ invariant. These results indicate that CsTi3Bi5 can host topological surface states within the local energy gaps, as reported in experimental studies [40, 38].
@@ III.2 VHSs and nematic instability
We now examine the kagome electronic features and their role in driving electronic nematic order in CsTi3Bi5. Figure 2(a) presents the Ti $d_{x^{2}-y^{2}}$/$d_{xy}$ orbital-resolved band structure, highlighting the Kagome electronic features near the Fermi level. The two VHSs, VHS1 and VHS2, are located at energies 0.22 eV and 0.74 eV above the Fermi level at $M$ point, respectively, while a flat band lies around -0.4 eV below it. These kagome bands cross the Fermi level, forming the Dirac cone at the $K$ point. To characterize the flavor of VHSs, we isolate the sublattice projections of the Bloch wave functions at VHSs, as shown in Fig.  2(b). Both VHS1 and VHS2 are sublattice pure, with each sublattice contributing VHS at different M points. Typically, in a kagome band structure, two VHSs exhibit a mix of sublattice pure and mixed characters; however, both VHSs here are sublattice pure without exhibiting mixed characters. This behavior arises from the three-dimensionality of VHS2, resulting from the mixing of Ti $d_{x^{2}-y^{2}}$/$d_{xy}$ with Ti $d_{z^{2}}$ orbitals. The sublattice-pure decoration of VHSs splits the three prominent particle-hole scattering channels in the hexagonal Fermi surface into six, thereby weakening the instabilities like CDW while promoting sublattice interference and long-range interactions [61] (see SM of additional details) [58].
To analyze the order of VHSs [62], we fit the energy dispersion around the saddle points (VHS1 and VHS2) along the $M-K-M$ direction using both quadratic and quartic polynomials, as shown in Fig. 2(c). The band dispersion is better described by a quartic polynomial $E=a+bk_{x}^{2}+ck_{x}^{4}$, with fitting parameters $a=0.22$, $b=1.07$, and $c=7.63$ for VHS1, and $a=0.74$, $b=-0.33$, and $c=-13.87$ for VHS2. The goodness-of-fit parameters, R-square and root mean square error (RMSE), are 0.9999 and 0.0014 for VHS1 and 0.9988 and 0.0003 for VHS2. These values, with R-squared close to 1 and RMSE near 0, confirm the excellent quality of the quartic polynomial fit. The deviation of energy dispersion from a normal quadratic form indicates the presence of high-order VHSs (Fig. 2(e)). The high-order VHSs warp the ideal hexagonal Fermi surface associated with normal VHSs, reducing Fermi surface nesting and associated instabilities. Simultaneously, they enhance the DOS, thereby amplifying correlations-driven many-body effects. The sublattice-pure high-order VHSs in CsTi3Bi5 thus suggest a decreased likelihood of CDW instabilities while enhancing sublattice interference and electron correlation effects [63, 64, 65]. This interplay can ultimately strengthen electronic instabilities. Such phenomena have been observed in materials like CsV3Sb5 [66] and high $T_{c}$ cuprates [67], where significant many-body effects prevail.
@@
The aforementioned discussion indicates the possibility of electronic nematic instability in CsTi3Bi5, as observed in scanning tunneling microscopy measurements [39]. Notably, the nematic order in CsTi3Bi5 reduces the $C_{6}$ rotation symmetry to $C_{2}$ symmetry. Due to the sublattice pure VHSs, a simple onsite potential can break rotation symmetry, distorting the Fermi surface and lowering the density of states at VHSs [30]. To test this phenomenology, we introduce an onsite potential in one of the sublattices, which effectively modifies the states in one of the three inequivalent $M$ points. The resulting $\chi({\bf q})$, shown in Fig. 2(h), demonstrates that the $C_{6}$ symmetry of the system is broken, leading to the emergence of $C_{2}$ symmetry. While this model captures the symmetry lowering in the absence of CDW ordering, the multi-orbital Fermi surface and sublattice interference result in non-uniform symmetry breaking, as seen in quasiparticle interference patterns, which warrant further investigation [39, 40, 41].
@@ III.3 Robustness of structural stability
We now examine the possibility of lattice instabilities in CsTi3Bi5. While the VHSs suggest possible electronic instabilities, they are situated slightly above the Fermi level. Therefore, it is essential to investigate dynamical instabilities that may arise from shifting the Fermi level to the VHSs or flat band energies. Figure 3(a) shows the phonon dispersion for pristine CsTi3Bi5, along with its electron and hole doping cases. In all cases, the phonon eigenvalues remain positive without any imaginary values across the full Brillouin zone. There are only minor changes in the phonon bands even under substantial doping of 1 electron (hole) per unit cell, which moves the Fermi level to the VHS1 (near flat band) energies (see SM) [58]. We also assessed the lattice stability by applying 1% compressive and tensile strains (Fig. 3(b)), which similarly yield stable phonon modes. The absence of imaginary phonon frequencies under various conditions suggests that CsTi3Bi5 is robust against lattice instabilities. Notably, even moving the Fermi level to the VHSs or flat bands through doping has no detrimental impact on lattice stability. These results align well with recent experimental reports indicating that CsTi3Bi5 exhibits no structural instabilities even when the VHSs are moved to the Fermi level through Cs doping [42].
@@
In bulk kagome materials, the kagome atomic layers are typically buried deep within the structure, complicating the assessment of their influence on structural instabilities. To examine the role of the Ti kagome layer, we calculated the surface phonon dispersions of CsTi3Bi5 for two surface terminations: one that directly exposes the Ti kagome layer and another that positions it just beneath the surface. The phonon spectra for both terminations, presented in Figs. 3(c) and 3(d), reveal no imaginary phonon frequencies in the surface Brillouin zone. These results indicate that CsTi3Bi5 retains structural stability even with the kagome layer is exposed at the surface, suggesting that any potential instabilities here would be driven by electronic factors. These characteristics position CsTi3Bi5 as a unique kagome system, especially in contrast to isostructural AV3Sb5 materials and bi-kagome layer systems such as AV6Sn6, which exhibit significant lattice instabilities in their surface-exposed kagome layers [68].